CSePub Home Book Store Instructors Independent Authors About A very careful new father wants to set money aside for his batby daughters wedding. If the wedding will take place in 21.00 years, he expects the wedding will cost $104,100.00. If the father can eam 11.00% APR with quarterly compounding on his investments, how much does he need to invest today to reach his goal? Answer Format: Currency: Round to: 2 decimal places. Enter Answer Here. Submit Answer -Prev Problem t Aa Problems «Wed Coppright 2019 Conley Sith Publishing All Right neserved Contact Us Pivacy Policy Tems
Answer #1

Amount after 21 years = 104,100
Rate Per Quarter = 11%/4
Number of Periods = 21*4 = 84

Amount to be paid today = 104,100/(1+11%/4)84 = 10,660.49

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well